# ContPhy: Continuum Physical Concept Learning and Reasoning from Videos

### v1

@@ 5 Experiments
Subset | Settings | Blind Models | Visual Models | MLLMs | ContPRO | Human | ||||||
RND | FRQ | B-LSTM | C-LSTM | MAC | HCRN | ALPRO | Gemini | GPT-4V | ||||
Rope | Prop. | 30.0 | 53.3 | 54.7 | 52.7 | 53.3 | 51.7 | 60.7 | 35.5 | 48.0 | 71.4 | 84.7 |
C-Opt. | 51.3 | 51.6 | 74.0 | 74.0 | 74.2 | 74.3 | 76.2 | 48.2 | 42.0 | 75.6 | 90.2 | |
C-Ques. | 14.7 | 19.0 | 46.0 | 45.0 | 39.8 | 48.1 | 50.7 | 12.0 | 11.3 | 48.8 | 75.0 | |
G-Opt. | 55.2 | 49.7 | 47.4 | 51.2 | 50.3 | 56.0 | 46.2 | 51.6 | 57.0 | 55.8 | 91.9 | |
G-Ques. | 4.5 | 11.2 | 7.9 | 6.7 | 6.7 | 2.3 | 1.1 | 10.3 | 12.1 | 2.3 | 84.0 | |
Fluid | Prop. | 33.3 | 52.7 | 49.3 | 54.0 | 30.0 | 52.7 | 48.0 | 10.0 | 25.0 | 78.0 | 75.8 |
C-Opt. | 52.9 | 57.9 | 56.1 | 55.0 | 56.5 | 52.6 | 56.8 | 47.3 | 53.3 | 75.7 | 82.5 | |
C-Ques. | 6.0 | 17.2 | 7.8 | 8.6 | 6.9 | 4.3 | 6.0 | 5.1 | 5.1 | 36.2 | 60.6 | |
G-Opt. | 59.9 | 63.1 | 57.3 | 57.3 | 51.2 | 67.7 | 62.7 | 44.4 | 53.8 | 77.3 | 75.0 | |
G-Ques. | 7.5 | 36.3 | 22.5 | 22.5 | 17.5 | 41.3 | 32.5 | 11.3 | 7.5 | 60.0 | 64.3 | |
P-Opt. | 53.8 | 50.1 | 51.4 | 51.4 | 53.5 | 50.6 | 53.8 | 52.4 | 50.0 | 90.1 | 73.9 | |
P-Ques. | 4.8 | 12.5 | 12.5 | 12.5 | 12.5 | 1.9 | 12.7 | 5.8 | 13.0 | 68.3 | 42.9 | |
Cloth | Prop. | 46.7 | 41.3 | 56.7 | 46.7 | 59.3 | 52.0 | 48.0 | 42.0 | 49.0 | 60.0 | 81.4 |
P-Opt. | 52.2 | 61.7 | 55.2 | 67.5 | 57.9 | 62.0 | 68.8 | 50.1 | 53.0 | 64.7 | 79.6 | |
P-Ques. | 46.0 | 56.7 | 42.3 | 57.3 | 50.7 | 56.3 | 57.3 | 43.0 | 47.5 | 60.0 | 77.3 | |
Ball | Prop. | 53.5 | 52.0 | 45.3 | 54.7 | 48.0 | 43.3 | 48.0 | 54.0 | 45.0 | 54.0 | 76.9 |
C-Opt. | 53.6 | 65.8 | 66.7 | 64.2 | 66.1 | 65.3 | 63.9 | 60.9 | 66.7 | 71.6 | 93.9 | |
C-Ques. | 30.4 | 48.7 | 43.4 | 41.8 | 3.3 | 28.7 | 40.2 | 29.6 | 46.9 | 57.4 | 90.9 | |
G-Opt. | 55.9 | 52.1 | 53.3 | 54.1 | 58.1 | 57.0 | 56.3 | 54.1 | 51.4 | 68.1 | 89.7 | |
G-Ques. | 30.2 | 38.5 | 16.7 | 20.0 | 18.9 | 38.9 | 4.4 | 24.6 | 18.0 | 52.2 | 84.6 | |
P-Opt. | 50.6 | 67.8 | 68.9 | 67.4 | 64.4 | 61.7 | 65.2 | 51.7 | 45.4 | 92.4 | 72.5 | |
P-Ques. | 25.9 | 51.7 | 45.5 | 45.5 | 46.6 | 1.1 | 3.4 | 25.9 | 17.2 | 88.6 | 58.8 |
In this section, we evaluate the ContPhy dataset. We first introduce the experimental setup and then analyze different models’ performance.
@@ 5.1 Experimental Setup
For simplicity, each physical property question is regarded as a classification task among all possible answers. Each dynamic question is treated as a binary classification task for each question-choice pair. For dynamic questions, we report the accuracy for each option and per question. A question is correct only if all choices in this multiple-choice question are correctly answered.
Blind Models. This family of models includes baselines that only rely on question input, to analyze language biases in ContPhy. RND chooses at random a possible answer, or randomly selects between true-false binary answer pairs for every multiple-choice question. FRQ selects the most frequent answer based on the question type. B-LSTM utilizes an LSTM (Hochreiter & Schmidhuber, 1997) to encode the questions only and predict answers.
Visual Models. These models incorporate both visual and language representations for answering questions. C-LSTM extracts video features via ResNet-50 convolutional neural network (CNN) (He et al., 2016) on 25 sampled frames of videos and averages them over time as the visual input. We concatenate this visual input with the question embedding from the last hidden state of LSTM to predict answers. HCRN (Le et al., 2020) uses conditional relational networks to learn relations hierarchically in the video, as well as the questions. MAC (Hudson & Manning, 2018) has competitive results on previous datasets, which uses co-attention mechanism to model both textual and visual information. ALPRO (Li et al., 2022a) is a popular model pre-trained on video-text corpus and achieved state-of-the-art results on several video-language datasets. We fine-tune ALPRO on our dataset based on the official pre-trained checkpoint.
@@ 5.2 Evaluation of Physical Reasoning
Performance of Blind Models. Blind models operate and respond solely to textual data, reflecting the quality and structure of question design. Generally, these models have weaker performance than other families of models, showing the importance of cooperating visual information to handle the questions in ContPhy. We also observe that blind models perform similarly to other model families in some scenarios like the goal-driven questions of the rope scenario. We think the reasons are that these dynamic questions are too challenging for current machine models, which require the understanding of physical commonsense and predict information that is not directly observable. Note that human beings can still easily achieve high performance.
@@
Performance of Visual Models. Visual models, which integrate both visual and language representations, exhibit relatively consistent performance across various questions. Among all dynamic questions, they excel on the rope’s counterfactual and cloth’s predictive, but fall short on goal-driven questions. The inherent complexity of goal-driven questions, which require reverse reasoning based on the goal, may account for its bad performance. Among different scenarios and visual models, ALPRO distinguishes itself by its robust overall performance, notably in cloth and rope, which shows the advantages of large-scale video-text pre-training and alignment, emphasizing its effectiveness in complex visual reasoning. Despite these advancements, no visual model has yet achieved top accuracy in all scenarios, underscoring the challenge and significance of our ContPhy.
@@
Performance of Multimodal Large Language Models. Compared with other baselines, both foundation models, Gemini and GPT4-V, fall short in cloth and fluid questions in both property and dynamics levels, showing that they fail to perceive highly deformable objects. However, they perform the best in certain question settings of rope and ball scenarios. Note that these models have never been trained on ContPhy and their visibility is limited to discrete frames in our setup. We think the reason is that objects in rope and ball, e.g. cubes and spheres, are more common to foundation models than those in fluid and cloth, e.g. different colored liquids and clothes. For example, GPT-4V shows its capabilities in counting objects and perceiving object colors, which probably raises the rope property evaluation score. MLLMs also distinguish themselves by their complex reasoning capabilities, accounting for high accuracy in rope’s goal-driven questions. We argue that current foundation models lack the necessary capability to infer physical properties and predict the dynamics of complex scenes with soft objects and fluid.
@@
Performance of ContPRO. ContPRO excels significantly over other machine models on most questions, particularly on property inference and predictive questions. Notably, its accuracy on predictive questions even surpasses human performance. A possible explanation is that we adopt MPM for fluid and ball scenarios, which can precisely predict objects’ positions in the short term, while humans tend to give vague estimations. Such an approach also accounts for its superior performance on the fluid’s goal-driven questions. In addition, ContPRO demonstrates exceptional performance on fluid property inference, achieving an accuracy of 96.9% on “stick number” and 63.5% on “density”. We believe the reason is that we have utilized a fine-tuned Mask R-CNN predictor to identify sticks in videos. For rope and ball scenarios, we employ DPI-Net, a GNN-based simulator, which cannot exhibit absolute advantages over visual models.
@@
Human Performance. We randomly sampled some video-question pairs from the test set to assess the human ability to comprehend the physical properties and dynamic events presented in both video and textual descriptions. To evaluate human performance on ContPhy, 16 people participated in the study. Participants were required to have fundamental English reading skills and a basic physical knowledge background. First, each participant was asked to select a scenario randomly, after which they were presented with distinct video-question pairs. Participants were instructed to answer with a phrase when presented with physical property questions, while for dynamics questions they were required to provide a binary true-false response from available choices. We obtained 460 valid human answers encompassing all scenarios and question types within ContPhy. We can observe from Table 2 that it beats visual models and foundation models in all scenarios. This shows the fundamental ability and strength of humans to perform visual reasoning and inference from videos.
@@
Evaluation Conclusion. The strong human results demonstrate that humans maintain a strong capacity to comprehend both videos and questions, make physical property inferences from given videos, and predict and reason counterfactual hypotheses concerning unseen information. Machine model results show that even state-of-the-art models struggle with answering these physical questions. This indicates that our dataset poses a significant challenge for vision-language models to achieve similar basic physical video understanding ability with human beings. We also propose an oracle model to demonstrate the potential to combine recent large language models with traditional particle-based dynamic simulation for effective physical reasoning.

### v2

@@ 5 Experiments
| Subset | Settings | Blind Models | Visual Models | Physical Models | MLLMs | ContPRO | Human | ||||||||
| RND | FRQ | B-LSTM | C-LSTM | MAC | HCRN | ALPRO | Violet | PhyDNet | PIP | Gemini | GPT-4V | ||||
| Rope | Prop. | 30.0 | 53.3 | 54.7 | 52.7 | 53.3 | 51.7 | 60.7 | 51.7 | 59.0 | 31.5 | 35.5 | 48.0 | 71.4 | 84.7 |
| C-Opt. | 51.3 | 51.6 | 74.0 | 74.0 | 74.2 | 74.3 | 76.2 | 76.0 | 77.7 | 75.2 | 48.2 | 42.0 | 75.6 | 90.2 | |
| C-Ques. | 14.7 | 19.0 | 46.0 | 45.0 | 39.8 | 48.1 | 50.7 | 43.1 | 47.9 | 48.3 | 12.0 | 11.3 | 48.8 | 75.0 | |
| G-Opt. | 55.2 | 49.7 | 47.4 | 51.2 | 50.3 | 56.0 | 46.2 | 55.2 | 54.4 | 50.6 | 51.6 | 57.0 | 55.8 | 91.9 | |
| G-Ques. | 4.5 | 11.2 | 7.9 | 6.7 | 6.7 | 2.3 | 1.1 | 1.1 | 5.6 | 2.2 | 10.3 | 12.1 | 2.3 | 84.0 | |
| Fluid | Prop. | 33.3 | 52.7 | 49.3 | 54.0 | 30.0 | 52.7 | 48.0 | 50.9 | 51.3 | 37.0 | 10.0 | 25.0 | 78.0 | 75.8 |
| C-Opt. | 52.9 | 57.9 | 56.1 | 55.0 | 56.5 | 52.6 | 56.8 | 60.4 | 59.5 | 49.1 | 47.3 | 53.3 | 75.7 | 82.5 | |
| C-Ques. | 6.0 | 17.2 | 7.8 | 8.6 | 6.9 | 4.3 | 6.0 | 1.7 | 10.3 | 6.0 | 5.1 | 5.1 | 36.2 | 60.6 | |
| G-Opt. | 59.9 | 63.1 | 57.3 | 57.3 | 51.2 | 67.7 | 62.7 | 67.3 | 55.9 | 67.7 | 44.4 | 53.8 | 77.3 | 75.0 | |
| G-Ques. | 7.5 | 36.3 | 22.5 | 22.5 | 17.5 | 41.3 | 32.5 | 41.2 | 40.0 | 41.3 | 11.3 | 7.5 | 60.0 | 64.3 | |
| P-Opt. | 53.8 | 50.1 | 51.4 | 51.4 | 53.5 | 50.6 | 53.8 | 53.2 | 51.7 | 45.5 | 52.4 | 50.0 | 90.1 | 73.9 | |
| P-Ques. | 4.8 | 12.5 | 12.5 | 12.5 | 12.5 | 1.9 | 12.7 | 3.8 | 4.8 | 3.8 | 5.8 | 13.0 | 68.3 | 42.9 | |
| Cloth | Prop. | 46.7 | 41.3 | 56.7 | 46.7 | 59.3 | 52.0 | 48.0 | 55.0 | 58.7 | 54.0 | 42.0 | 49.0 | 60.0 | 81.4 |
| P-Opt. | 52.2 | 61.7 | 55.2 | 67.5 | 57.9 | 62.0 | 68.8 | 68.2 | 63.5 | 61.6 | 50.1 | 53.0 | 64.7 | 79.6 | |
| P-Ques. | 46.0 | 56.7 | 42.3 | 57.3 | 50.7 | 56.3 | 57.3 | 55.7 | 47.3 | 46.3 | 43.0 | 47.5 | 60.0 | 77.3 | |
| Ball | Prop. | 53.5 | 52.0 | 45.3 | 54.7 | 48.0 | 43.3 | 48.0 | 48.0 | 52.7 | 54.0 | 54.0 | 45.0 | 54.0 | 76.9 |
| C-Opt. | 53.6 | 65.8 | 66.7 | 64.2 | 66.1 | 65.3 | 63.9 | 65.6 | 67.2 | 63.7 | 60.9 | 66.7 | 71.6 | 93.9 | |
| C-Ques. | 30.4 | 48.7 | 43.4 | 41.8 | 3.3 | 28.7 | 40.2 | 41.8 | 44.3 | 24.6 | 29.6 | 46.9 | 57.4 | 90.9 | |
| G-Opt. | 55.9 | 52.1 | 53.3 | 54.1 | 58.1 | 57.0 | 56.3 | 57.4 | 57.4 | 54.1 | 54.1 | 51.4 | 68.1 | 89.7 | |
| G-Ques. | 30.2 | 38.5 | 16.7 | 20.0 | 18.9 | 38.9 | 4.4 | 21.1 | 21.1 | 22.2 | 24.6 | 18.0 | 52.2 | 84.6 | |
| P-Opt. | 50.6 | 67.8 | 68.9 | 67.4 | 64.4 | 61.7 | 65.2 | 64.4 | 67.4 | 62.9 | 51.7 | 45.4 | 92.4 | 72.5 | |
| P-Ques. | 25.9 | 51.7 | 45.5 | 45.5 | 46.6 | 1.1 | 3.4 | 2.3 | 17.0 | 6.8 | 25.9 | 17.2 | 88.6 | 58.8 | |
In this section, we evaluate the ContPhy dataset. We first introduce the experimental setup and then analyze the performance of different models.
@@ 5.1 Experimental Setup
For simplicity, each physical property question is regarded as a classification task among all possible answers. Each dynamic question is treated as a binary classification task for each question-choice pair. For dynamic questions, we report the accuracy of per option per question. A question is correct only if all choices in this multiple-choice question are correctly answered.
Blind Models.
This family of models includes baselines that only rely on question input, to analyze language biases in ContPhy. RND chooses at random a possible answer, or randomly selects between true-false binary answer pairs for every multiple-choice question. FRQ selects the most frequent answer based on the question type. B-LSTM utilizes an LSTM (Hochreiter & Schmidhuber, 1997) to encode the questions only and predict answers.
Visual Models.
These models incorporate both visual and language representations for answering questions. C-LSTM extracts video features via ResNet-50 convolutional neural network (CNN) (He et al., 2016) on 25 sampled frames of videos and averages them over time as the visual input. We concatenate this visual input with the question embedding from the last hidden state of LSTM to predict answers. HCRN (Le et al., 2020) uses conditional relational networks to learn relations hierarchically in the video, as well as the questions. MAC (Hudson & Manning, 2018) has competitive results on previous datasets, which uses a co-attention mechanism to model both textual and visual information. ALPRO (Li et al., 2022a) is a popular model pre-trained on video-text corpus and achieved state-of-the-art results on several video-language datasets. We fine-tune ALPRO on our dataset based on the official pre-trained checkpoint.
Physical Models.
These specialized models are trained for physical reasoning. PhyDNet (Le Guen & Thome, 2020) is a two-branch deep architecture, which explicitly disentangles PDE dynamics from unknown complementary information. PIP (Duan et al., 2021) utilizes a deep generative model to model mental simulations, in order to predict future physical interactions. To evaluate our benchmark on these physical baselines, we first generated object masks based on each question and fed them into models, together with the video features. For the open-ended questions, we added a fully connected layer to predict the answer labels with a cross-entropy loss.
Multimodal Large Language Models (MLLMs).
@@ 5.2 Evaluation of Physical Reasoning
Blind models operate and respond solely to textual data, reflecting the quality and structure of question design. Generally, these models have weaker performance than other families of models, showing the importance of cooperating visual information to handle the questions in ContPhy. We also observe that blind models perform similarly to other model families in some scenarios like the goal-driven questions of the rope scenario. We think the reasons are that these dynamic questions are too challenging for current machine models, which require the understanding of physical commonsense and predict information that is not directly observable. Note that human beings can still easily achieve high performance.
Performance of Visual Models.
@@
Visual models, which integrate both visual and language representations, exhibit relatively consistent performance across various questions. Among all dynamic questions, they excel on the rope’s counterfactual and cloth’s predictive, but fall short on goal-driven questions. The inherent complexity of goal-driven questions, which require reverse reasoning based on the goal, may account for its bad performance. Among different scenarios and visual models, ALPRO distinguishes itself by its robust overall performance, notably in cloth and rope, which shows the advantages of large-scale video-text pre-training and alignment, emphasizing its effectiveness in complex visual reasoning. Despite these advancements, no visual model has yet achieved top accuracy in all scenarios, underscoring the challenge and significance of our ContPhy.
Performance of Physical Models.
Physical models are specialized models designed for special physical tasks. These models achieve competitive overall performance and excel in some settings, such as PhyDNet on rope’s and ball’s counterfactual, and PIP on fluid’s goal-driven. However, these specialized models also have limitations in some scenarios, such as cloth. We hypothesize the reasons are that these models are mainly designed for physical reasoning tasks with simple visual primitives, like sphere collision and movement. However, our dataset focuses on continuum objects in diverse environments and different question types, which makes it difficult for these models to grasp the physical rules behind the scenarios.
Performance of Multimodal Large Language Models.
@@
Compared with other baselines, both foundation models, Gemini and GPT4-V, fall short in cloth and fluid questions in both property and dynamics levels, showing that they fail to perceive highly deformable objects. However, they perform the best in certain question settings of rope and ball scenarios. Note that these models have never been trained on ContPhy and their visibility is limited to discrete frames in our setup. We think the reason is that objects in rope and ball, e.g. cubes and spheres, are more common to foundation models than those in fluid and cloth, e.g. different colored liquids and clothes. For example, GPT-4V shows its capabilities in counting objects and perceiving object colors, which probably raises the rope property evaluation score. MLLMs also distinguish themselves by their complex reasoning capabilities, accounting for high accuracy in rope’s goal-driven questions. We argue that current foundation models lack the necessary capability to infer physical properties and predict the dynamics of complex scenes with soft objects and fluid.
Performance of ContPRO.
@@
ContPRO excels significantly over other machine models on most questions, particularly on property inference and predictive questions. Notably, its accuracy on predictive questions even surpasses human performance. A possible explanation is that we adopt MPM for fluid and ball scenarios, which can precisely predict objects’ positions in the short term, while humans tend to give vague estimations. Such an approach also accounts for its superior performance on the fluid’s goal-driven questions. In addition, ContPRO demonstrates exceptional performance on fluid property inference, achieving an accuracy of 96.9% on “stick number” and 63.5% on “density”. We believe the reason is that we have utilized a fine-tuned Mask R-CNN predictor to identify sticks in videos. For rope and ball scenarios, we employ DPI-Net, a GNN-based simulator, which cannot exhibit absolute advantages over visual models.
Human Performance.
@@
We randomly sampled some video-question pairs from the test set to assess the human ability to comprehend the physical properties and dynamic events presented in both video and textual descriptions. To evaluate human performance on ContPhy, 16 people participated in the study. Participants were required to have fundamental English reading skills and a basic physical knowledge background. First, each participant was asked to select a scenario randomly, after which they were presented with distinct video-question pairs. Participants were instructed to answer with a phrase when presented with physical property questions, while for dynamics questions they were required to provide a binary true-false response from available choices. We obtained 460 valid human answers encompassing all scenarios and question types within ContPhy. We can observe from Table 2 that it beats visual models and foundation models in all scenarios. This shows the fundamental ability and strength of humans to perform visual reasoning and inference from videos.
Evaluation Conclusion.
@@
The strong human results demonstrate that humans maintain a strong capacity to comprehend both videos and questions, make physical property inferences from given videos, and predict and reason counterfactual hypotheses concerning unseen information. Machine model results show that even state-of-the-art models struggle with answering these physical questions. This indicates that our dataset poses a significant challenge for vision-language models to achieve similar basic physical video understanding ability with human beings. We also propose an oracle model to demonstrate the potential to combine recent large language models with traditional particle-based dynamic simulation for effective physical reasoning.
6 Limitations
Our proposed benchmark, ContPhy, aims to complement existing physical reasoning benchmarks by encompassing diverse physical property inference across various scenarios and predicting corresponding dynamics. However, ContPhy still has limitations.
Language Diversity. While the synthesized questions generated by the question engine can effectively test AI models’ physical reasoning capabilities across diverse scenarios involving different objects, the language diversity remains limited. The current set of questions relies on a predefined vocabulary, resulting in a gap compared to natural language.
Scenario Complexity. We have carefully designed four distinct scenarios featuring various objects (e.g., solids, ropes, clothes, and fluids). However, real-world physical interactions can be considerably more complex, involving additional objects and physical factors not currently included in the dataset.
@@ 7 Conclusion
Acknowledgement
This work was supported by DSO grant DSOCO21072. We would also like to thank the computation support from AiMOS, a server cluster for the IBM Research AI Hardware Center.
Impact Statement
This paper presents work whose goal is to advance the field of Physical Commonsense Reasoning. We believe our work is useful for 1) evaluating the performance of current existing machine learning models for physical reasoning, and 2) facilitating researchers to develop more powerful AI models with physical commonsense. There are no potential negative societal consequences of our work that we feel must be specifically highlighted here.